How to find sin120

Answer by Boreal(15235)

sin 120=sin (90+30)
sin (a+b)=sin a* cos b+cos a*sin b
This is sin 90*cos 30+ cos 90*sin 30; cos 90=0, so second term is 0.
sin 90 is 1, so sin 120=cos 30=sqrt(3)/2